In Problems 7 and 8 find the general solution of the given differential equation.

8. y′′ + 2y′ + 5y = g(t),

(a) g(t) = −2t + 4t2;

(b) g(t) = t3;
Answer #1

o co solve this clifferential equation by method of uncentermined coefficient consider given differential equation y +2y! +5Al! we guess that yo= A +@4C-22 is pesticuler solution . Now we have to detemine the values of AB, AC :consider, ya AtBt+c+2Yct= SACH f 9 pct. YCH - ēti a casettasinzt) + Bolution of O. 22-2671 vie is generel Of considered given differential equatioii 65c+60=0 2°C--6G!). The locale *** *5872260-- BJECD-ke) = =+60+he) B=+(60+)* 4)) 3 ($ ) - :: 5 Af 2B+20=0 5 5A=-23-20. a A